Solve the system of equations by the Gaussian elimination method. (If the system is dependent, enter a general solution in terms of c1, c2, ...
3x − 5y + 2z = 2
x − 3y + 2z = 2
5x − 11y + 6z = 6

Answer by ewatrrr(24785)

 
Hi,
3x − 5y + 2z = 2
x − 3y + 2z = 2
5x − 11y + 6z = 6
Dependent system as D =0, as well as Dx=0, Dy= 0 and Dz = 0
or as demonstated by selecting different pairs of EQs and eliminating z for both:
2x - 2y = 0
2x - 2y = 0 |Subtracting 2nd from 1st
0 = 0 Dependent System
Dependent relationship in terms of z: (z-1, z-1, z)
x = z - 1 |Eliminating y-variable using a pair of EQs
y = z - 1 |Eliminating x-variable using a pair of EQs